# Multilingual large language models leak human stereotypes across language boundaries

### v2

@@ Abstract.
Multilingual large language models have been increasingly popular for their proficiency in processing and generating text across various languages. Previous research has shown that the presence of stereotypes and biases in monolingual large language models can be attributed to the nature of their training data, which is collected from humans and reflects societal biases. Multilingual language models undergo the same training procedure as monolingual ones, albeit with training data sourced from various languages. This raises the question: do stereotypes present in one social context leak across languages within the model? In our work, we first define the term “stereotype leakage” and propose a framework for its measurement. With this framework, we investigate how stereotypical associations leak across four languages: English, Russian, Chinese, and Hindi. To quantify the stereotype leakage, we employ an approach from social psychology, measuring stereotypes via group-trait associations. We evaluate human stereotypes and stereotypical associations manifested in multilingual large language models such as mBERT, mT5, and GPT-3.5. Our findings show a noticeable leakage of positive, negative, and non-polar associations across all languages. Notably, Hindi within multilingual models appears to be the most susceptible to influence from other languages, while Chinese is the least. Additionally, GPT-3.5 exhibits a better alignment with human scores than other models.
@@ 1. Introduction
Cultural stereotypes about social groups can be transmitted based on how these social groups are represented, treated, and discussed within each culture (Martinez et al., 2021; Lamer et al., 2022; Rhodes et al., 2012). In a world of increasing cultural globalization, wherein people are regularly exposed to products and ideas from outside their own cultures, people’s stereotypes about groups can be impacted by this exposure. For instance, blackface is characterized as one of America’s first cultural exports, as the performance of American minstrelsy shows in different countries popularized racist depictions of Black Americans within those other cultures (Thelwell, 2020).
Recently, the deployment of large language models has the potential to exacerbate the issue. Large language models are becoming increasingly language-agnostic. For instance, models like ChatGPT111https://openai.com/chatgpt and mBART (Lin et al., 2022) can operate without being restricted to a specific language, handling input and output in multiple languages simultaneously. This thus gives rising opportunities for what we refer to as stereotype leakage, or the transmission of stereotypes from one culture to another.
Stereotypes are abstract and over-generalized pictures drawn about people based on their group membership, and these perceptions can be specific to each culture. Stereotype leakage within large language models may export harmful stereotypes across cultures and reinforce Anglocentricism222Anglocentrism is the practice of viewing and interpreting the world from an English-speaking perspective with the prioritization of English culture, language, and values. Anglocentrism can lead to biases and neglect of global perspectives and experiences.. Previous works (e.g., Goldstein et al., 2023; Weidinger et al., 2021) have highlighted the potential for language model outputs to change users’ perceptions and behaviors. Stereotype leakage from large language models may further entrench existing stereotypes among model users, as well as create new stereotypes that have transferred from a different language. Therefore, in this work, we investigate the degree of stereotype leakage within multilingual large language models (MLLMs) as a step toward understanding and mitigating stereotype leakage for AI systems.
Large language models are currently the backbone of many natural language processing (NLP) models. MLLMs are language models pre-trained with a large amount of data from multiple languages so that they can process NLP tasks in various languages as well as cross-lingual tasks. Recent MLLMs, such as GPT models (Brown et al., 2020; OpenAI, 2023) designed for standalone applications and models such as mBERT (Müller et al., 2020), XLM (Lample and Conneau, 2019), mT5 (Xue et al., 2020), mBART (Lin et al., 2022), intended for use as back-end tools, show satisfactory performance on NLP tasks across around 100 languages. One major advantage of such models is that low-resource languages (languages with less training data) can benefit from high-resource languages through shared vocabulary (Lample and Conneau, 2019) and structural similarities (word-ordering or word-frequency) (K et al., 2020).
@@
Large language models are trained on existing language data, and even monolingual language models have been demonstrated to replicate stereotypical associations present in the training data. (Nadeem et al., 2020; Nangia et al., 2020a; Cao et al., 2022). Thus, with the shared knowledge between languages in MLLMs, it is likely that stereotypes may also leak between languages. Though LLMs are trained on language-based data rather than culture-based data, languages reflect the stereotypes associated with the cultures they represent. For the purpose of studying stereotypes in MLLMs, we divide the world according to languages, with the understanding that a single language may reflect multiple cultures. Previously, many works have examined Western stereotypes in English language models (e.g. Nadeem et al., 2020; Nangia et al., 2020a; Cao et al., 2022), whereas limited works have attempted to assess stereotypes in multilingual language models (e.g. Kaneko et al., 2022; Levy et al., 2023; Câmara et al., 2022) due to the complexity of stereotypes manifested in various cultures, limited resources, and Anglocentric norms (Talat et al., 2022).
@@
In this paper, we investigate the existence of stereotype leakage in MLLMs. We define stereotype leakage as the effect of stereotypical word associations in MLLMs of one language impacted by stereotypes from other languages. We conduct a human study to collect human stereotypes, adopt word association measurement approaches from previous works (Cao et al., 2022; Kurita et al., 2019) to measure stereotypical associations in MLLMs, and analyze the strength and nature of stereotype leakage across different languages both quantitatively and qualitatively.
To test our hypothesis that there is significant stereotype leakage across languages in MLLMs, we sample four languages: English, Russian, Chinese, and Hindi. We pick languages that come from the Indo-European and Sino-Tibetan language families, ranging from high (English) to low-resource (Hindi) languages333 High-resource languages are languages that have more training data available, while low-resource languages have less.. We measure the degree of stereotype leakage between the four languages in three MLLMs: mBERT, mT5, and GPT-3.5444Both the code and the dataset, along with a datasheet (Gebru et al., 2018), are available under a MIT licence at: https://github.com/AnnaSou/Stereotype_Leakage.. Both mBERT and mT5 are back-end MLLMs. MT5 has better multilingual performance than mBERT, whereas mBERT has more comparable monolingual BERT models for the four languages. GPT-3.5 is one of the state-of-the-art MLLMs that has been popularly deployed to users. With these, we examine the impact of human stereotypes from different languages on stereotypical associations in MLLMs.
2. Measuring Stereotype Leakage in MLLMs
For each language, we aim to assess the degree of stereotype leakage from the other languages to this target language in MLLMs. Specifically, we measure the effect of human stereotypes from all four languages () on the target language’s MLLM stereotypical association (), as shown in Equation 1. We also control the impact of the stereotypical association from the target monolingual model (). However, since we can only find monolingual BERT model for all four languages, we use these as proxies of for all MLLMs. We use a mixed-effect model to fit the formula and calculate the effect. If the coefficient of a variable is positive and has a p-value of less than , then the variable has a significant effect on . If there are significant effects from a non-target language’s human stereotypes, then there is potential stereotype leakage from this non-target language to the target language.
In the following section, we discuss how we measured each of the variables.
2.1. Stereotype Measurement
In this paper, we measure stereotypes through group-trait associations with traits from the Agency Beliefs Communion (ABC) model of stereotype content (Koch et al., 2020). The model consists of 16 trait pairs (each pair represents two polarities) that are designed to characterize group stereotypes along the dimensions of agency/socioeconomic success, conservative–progressive beliefs, and communion, as listed in Table 1.
If a group (e.g. “immigrant”, “Asian person”) has a high degree of association with a trait (e.g. religious, confident), then we consider that trait a stereotype of the group. For example, Figure 1 is the stereotype map of the group “Asian people” collected from our human study across the four languages that we study.
For the groups, we picked groups listed in Table 2: shared groups with shared stereotypes (groups that are present in all four countries and are expected to be targeted by similar stereotypes), shared groups with non-shared stereotypes (groups that are present in all four countries but expected to be targeted by dissimilar stereotypes), and non-shared groups (groups that exist uniquely in each country). For shared groups, we manually selected groups from the list of social groups from (Cao et al., 2022). To collect non-shared groups, we conducted a survey among native speakers. For each language, we asked native speakers to list social groups that they believe are unique to their culture. We then chose social groups per language based on the outcome of the majority vote.
@@ 2.1. Stereotype Measurement
| low status | high status | religious | science-oriented | dishonest | sincere | |||||
| dominated | dominating | conventional | alternative | cold | warm | |||||
| poor | wealthy | conservative | liberal | benevolent | threatening | |||||
@@
| Non-shared/ Non-shared | USA: Texan, Mormon, Puerto Rican |
| Russia: VDV soldier, Muscovite, Chechenets | |
| China: migrant worker, Hui person, Shanghainese person | |
| India: Brahmin person, Gujarati person, Shudra person |
2.1.1. Human stereotypes
@@
To collect human stereotypes, we conduct a human study on Prolific555https://www.prolific.co/ for each of the four languages with native speakers of the respective languages who lived or still live in the United States, Russia, China, and India666Approved by our institutional IRB, #1724519-3.. In the survey, participants are first asked to mark at least social groups that they feel they are familiar with. Then they are asked to rate the group-trait associations of social groups from their list of familiar groups. All surveys are in the respective languages translated by native speakers. For shared/shared and shared/non-shared groups, we collect at least participants’ annotations per group per language. For non-shared groups with non-shared stereotypes, we collect at least annotations for the language they originate from, with no minimum limit of annotations for other languages. See Appendix A for the survey details and participant demographics.
2.1.2. Model stereotypical associations
@@
To measure stereotypical group-trait associations in large language models, we adopt different approaches for different MLLMs. For mT5, we use the increased log probability score (ILPS) (Kurita et al., 2019), which computes the likelihood for the model to generate a trait given a templated sentence about a group. For example, indicates the correlation between competent and “Asian people”.
@@
For mBERT, we use the sensitivity test (SeT) (Cao et al., 2022), which is shown to have better alignment with human stereotypes (Cao et al., 2022). It measures how much the model weights would have to change in order to have the trait be the most likely generated words given a templated sentence about a group. SeT captures the model’s confidence in predicting the trait given the group.
In evaluating GPT-3.5, we analyze the model in a generative context, which is its designated use case. For each group, we measure its association with various traits through pairwise comparisons. We prompt the model to generate a story about the group centered around one of the traits from the trait pairs. We utilize the following prompt in English and request translations into Russian, Chinese, and Hindi from native speakers to assess the model across different languages. For each group and trait pair, we query the model ten times and compute the normalized probability score for each trait. Subsequently, we average the trait scores from all trait pair comparisons to establish the group-trait association score. Our experimentation employs the gpt-3.5-turbo-0125 model. Notably, we observe a tendency in the GPT model to favor positive traits (e.g., powerful, trustworthy) over negative traits (e.g., powerless, untrustworthy). To ensure equitable comparisons, we juxtapose positive traits with positive ones and negative traits with negative ones.
prompt sys = “You are a writer.”
prompt = “You are writing about [GROUP]. Before writing, think about what theme you want to pick. You can choose either ‘[TRAIT1]’ or ‘[TRAIT2]’ as your theme. You can also choose ‘neither’ if you think neither of these themes fits. Note that you can choose only one theme. Output the exact name of the theme only, without any punctuation.”
When processing the outputs of GPT-3.5, we employ an exact match criterion to assign scores to traits. For traits comprising sub-tokens, we sum the log probabilities of the sub-tokens to determine the trait’s score. In instances where outputs do not precisely match the traits in the prompt, we have them manually processed by native speakers of the respective language. Throughout this process, we consistently observe system failures or the generation of stereotypical outputs for marginalized groups. Notably, for certain groups like “feminist” and “Muslim person” in Chinese, the model often disregards the prompt and simply outputs the group name. Moreover, in some cases, the model alters the trait specified in the prompt. For example, it changes dominating to dominated for “disabled person” in English or poor to wealthy for “migrant worker” in Russian. Additionally, the model may overlook the traits provided in the prompt and generate stereotypical traits instead. For instance, in Russian, it generates rape and patriot for “Puerto Rican” or cowboy for “Texan”. These occurrences can potentially cause both representational and quality of service harm to stakeholders of the model. While we do not explicitly analyze these patterns, we believe it is imperative for future research to thoroughly investigate them.
3. Stereotype Leakage and Its Effects
In this section, we present our quantitative and qualitative results of the assessment of stereotype leakage across languages in MLLMs. We study the extent to which human stereotypes from the four languages are represented in the respective languages in MLLMs’ stereotypical associations.
3.1. Quantitative Results
In our analysis of mBERT, we observe significant leakages of stereotypes from Hindi to English and Chinese with coefficients of () and (), respectively. We also observe English human stereotypes manifesting in mBERT Hindi with a coefficient of (). Within the mT5 model, we find two significant stereotype leakages, both of which are leakages targeting Hindi. Russian and Chinese human stereotypes manifest in mT5 Hindi with coefficients of () and (), respectively. For GPT-3.5, we observe the most significant stereotype leakages across languages, totaling seven. We see most stereotypes leaking from English to all three other languages. The largest flows are from English to Chinese and Hindi, with coefficients of (). Meanwhile, all languages are prone to be affected by leakages from other languages. Overall, we observe that GPT-3.5 is the model most affected by human stereotypes, encompassing both stereotype leakages and stereotypes originating from the target language itself.
@@ 3.1. Quantitative Results
| Monolingual BERT | EN | RU | ZH | HI |

### v3

@@ Abstract.
Multilingual large language models have gained prominence for their proficiency in processing and generating text across languages. Like their monolingual counterparts, multilingual models are likely to pick up on stereotypes and other social biases present in their training data. In this paper, we study a phenomenon we term “stereotype leakage”, which refers to how training a model multilingually may lead to stereotypes expressed in one language showing up in the models’ behavior in another. We propose a measurement framework for stereotype leakage and investigate its effect across English, Russian, Chinese, and Hindi and with GPT-3.5, mT5, and mBERT. Our findings show a noticeable leakage of positive, negative, and non-polar associations across all languages. We find that of these models, GPT-3.5 exhibits the most stereotype leakage, and Hindi is the most susceptible to leakage effects.
@@ 1. Introduction
Large language models (LLMs) are trained on existing language data, and monolingual language models have been demonstrated to replicate stereotypical associations present in the training data. (Nadeem et al., 2020; Nangia et al., 2020; Cao et al., 2022). Multilingual large language models (MLLMs) are language models pre-trained with a large amount of data from multiple languages so that they can process NLP tasks in various languages as well as cross-lingual tasks. Previously, many studies have examined Western stereotypes in English language models (e.g. Nadeem et al., 2020; Nangia et al., 2020; Cao et al., 2022), whereas limited work has attempted to assess stereotypes in multilingual language models (e.g. Kaneko et al., 2022; Levy et al., 2023; Câmara et al., 2022) due to the complexity of stereotypes manifested in various cultures, limited resources, and Anglocentric norms (Talat et al., 2022). However, the study of stereotypes in multilingual settings is more complex than examining them within each language. Due to the shared knowledge between languages in MLLMs, it is likely that stereotypes expressed in one language are transmitted to the models’ behavior in another language.
@@
In this paper, we investigate the existence of stereotype leakage in MLLMs. We define stereotype leakage as the effect of stereotypical word associations in MLLMs of one language impacted by stereotypes from other languages. We conduct a human study to collect human stereotypes, adopt word association measurement approaches from previous studies (Cao et al., 2022; Kurita et al., 2019) to measure stereotypical associations in MLLMs and analyze the strength and nature of stereotype leakage across different languages both quantitatively and qualitatively.
Recent advancements in MLLMs have made them increasingly language-agnostic. For instance, models from GPT-family and mBART (Lin et al., 2022) can operate without being restricted to a specific language, simultaneously handling input and output in multiple languages. This thus gives rising opportunities for what we refer to as stereotype leakage from one culture to another.111Although language models are trained on language-based data rather than culture-based data, languages inherently reflect the stereotypes associated with their respective cultures. To study stereotypes in MLLMs, we divide the world by languages, recognizing that a single language may represent multiple cultures. Cultural stereotypes about social groups are shaped based on how these social groups are represented, treated, and discussed within each culture (Martinez et al., 2021; Lamer et al., 2022; Rhodes et al., 2012). Hence, people’s stereotypes about groups can be impacted by exposure to products and ideas from outside their own cultures. MLLMs, being the backbone of many natural language processing (NLP) applications, have the potential to exacerbate this issue by exporting harmful stereotypes across cultures and reinforcing Anglocentrism (Talat et al., 2022; joshi-etal-2020-state).222Anglocentrism is the practice of viewing and interpreting the world from an English-speaking perspective with the prioritization of English culture, language, and values. Anglocentrism can lead to biases and neglect of global perspectives and experiences.
We investigate the degree of stereotype leakage in MLLMs as a step toward understanding and mitigating this issue in AI systems. We test our hypothesis of significant stereotype leakage across languages in MLLMs by sampling four languages: English, Russian, Chinese, and Hindi. We choose languages from different writing systems—Latin alphabet, Cyrillic alphabet, Chinese characters, and Devanagari script—to enable a comprehensive evaluation of stereotype leakages in MLLMs. The models we assess are mBERT, mT5, and GPT-3.5. Based on our findings, all models demonstrate varying degrees of stereotype leakage, which occurs bidirectionally across languages without a dominant directionality. Among the models tested, GPT-3.5 exhibits the highest degree of stereotype leakage. Importantly, the stereotype leakage includes not only negative stereotypes but also positive and non-polar associations. Furthermore, our study reveals that stereotypes in other languages about social groups unfamiliar to those cultures are shaped by the stereotypes present in the native language. This indicates that multilingual language models reflect and propagate cultural stereotypes across linguistic boundaries.
@@
| low status | high status | religious | non-religious | dishonest | sincere | |||||
| dominated | dominating | irrational | rational | cold | warm | |||||
| poor | wealthy | conservative | liberal | threatening | benevolent | |||||
@@
2. Background and Related Work
Assessing multi-cultural biases and stereotypes in multilingual settings is challenging. As noted by Talat et al. (2022), there is a significant lack of benchmark datasets for measuring multilingual fairness. While many datasets exist in English, simply translating these datasets poses issues due to linguistic and cultural disparities. Furthermore, many existing fairness evaluation datasets are rooted in Western cultures, resulting in a gap that fails to encompass global cultural perspectives. Bartl et al. (2020) also highlighted the difficulty of measuring gender biases in languages with rich morphology and gender marking.
Many studies thus have been devoted to expanding the language boundary to assess the presence and impact of biases in multilingual settings by proposing new measurement approaches and evaluation datasets. Wang et al. (2021) focused on evaluating the multilingual fairness of pre-trained multimodal representations. Many studies delve deeply into gender biases in multilingual settings. Zhao et al. (2020) focused on word representations, while both Kaneko et al. (2022) and Steinborn et al. (2022) investigated gender bias in masked language models, each proposing new datasets for analyses. Furthermore, Touileb et al. (2022) examined occupational biases within Norwegian and multilingual language models, seeking to identify and mitigate these biases. Addressing intersectional biases, Câmara et al. (2022) mapped biases in sentiment analysis systems across English, Spanish, and Arabic, proposing a framework to measure these biases effectively. Additionally, Névéol et al. (2022) extended the CrowS dataset (Nangia et al., 2020) of sentence pairs in English for measuring bias in masked language models to the French language.
Going further, Dev et al. (2023) emphasized cultural inclusiveness by developing a stereotype dataset centered on Indian culture. This work highlights the importance of capturing local cultural contexts through community engagement. On the other hand, Levy et al. (2023) compared biases arising from multilingual training. Their findings show that biases are influenced by cultural contexts and often amplified during multilingual fine-tuning, underscoring the complexities involved in achieving fairness in multilingual NLP systems.
3. Measuring Stereotype Leakage in MLLMs
In measuring stereotype leakage in MLLMs, we aim to evaluate the extent to which stereotypes present in one language (the source language of stereotype leakage) become manifested in the models’ behavior in another language (the target language of stereotype leakage) as a consequence of multilingual training. Specifically, we measure the effect of stereotypical word associations in MLLMs of the target language impacted by stereotypes from source languages. These stereotypical associations reflect stereotypes manifested in the models’ behavior. In multilingual training, these can originate from stereotypes of the target language itself or from other languages. Meanwhile, some of these stereotypes are also captured during monolingual training. In this study, we aim to measure the effect of stereotype leakage from other languages as a result of multilingual training.
To do so, we form Equation 1, where is the stereotypical word associations in MLLMs of the target language and are human stereotypes of source languages. We measure the effect of stereotype leakage from source languages () to the target language of the MLLMs (). In this study, we sample four languages to study stereotype leakage: English (EN), Russian (RU), Chinese (ZH), and Hindi (HI). We choose languages that do not have shared orthography to focus on non-trivial leakages. These four languages come from the Indo-European and Sino-Tibetan language families, ranging from high (English) to low-resource (Hindi) languages.
and are all dimensional matrices, where is the number of traits and is the number of social groups. Each entry in the matrices is the stereotypical association score for a specific trait of a social group. is the intercept.
To measure the effect of stereotypes captured only due to multilingual training, we separate the stereotypes that may be captured during monolingual training by adding the variable, which is the stereotypical association from the target language’s monolingual model. Similarly, is also a dimensional matrix. Note that since we can only find the monolingual BERT model for all four languages, we use these as proxies of for all MLLMs.333The monolingual BERT models we use are bert- base-cased, bert-base-chinese, cointegrated/rubert-tiny2, and neuralspace- reverie/indic-transformers-hi-bert.
With the formula, the goal is to estimate how H of each language affects the . We use a mixed-effect model to fit the formula and calculate the effect. The mixed effect model fits a linear regression to the formula with the data. In our case, we set traits as the random effect variable. With the fitted linear model, we get the coefficients ( and ). The value indicates how much the stereotype score would change when the score changes. If the coefficient ( and ) of a variable is positive and has a p-value of less than , then we know that the variable has a significant effect on . If there are significant effects from a non-target language’s human stereotypes to , then there is potential stereotype leakage from this non-target language to the target language.
@@ 3. Measuring Stereotype Leakage in MLLMs
| Non-shared/ Non-shared | USA: Texan, Mormon, Puerto Rican Russia: VDV soldier, Muscovite, Chechenets China: migrant worker, Hui person, Shanghainese person India: Brahmin person, Gujarati person, Shudra person |
In our experiments, we focus on mBERT, mT5, and GPT-3.5.444The multilingual large language model versions we use are bert-base-multilingual-cased (mbert), google/mt5-base (mt5), and gpt-3.5-turbo-0125 version (Ouyang et al., 2022). Both mBERT and mT5 are back-end MLLMs. MT5 has better multilingual performance than mBERT, whereas mBERT has more comparable monolingual BERT models for the four languages. GPT-3.5 is one of the state-of-the-art MLLMs that has been popularly deployed to users.
With these, we examine the effect of stereotype leakages in MLLMs.555The code and the dataset, along with a datasheet (Gebru et al., 2018), will be published on GitHub. In the following section, we discuss how we measure each of the variables.
3.1. Stereotype Measurement
In this paper, we adopt a theory-grounded stereotype measurement approach from Cao et al. (2022). We measure stereotypes through group-trait associations with traits from the Agency Beliefs Communion (ABC) model of stereotype content (Koch et al., 2020), a well-established stereotype model constructed through social psychology studies for assessing human stereotypes. The model consists of 16 trait pairs (each pair represents two polarities) that are designed to characterize group stereotypes along the dimensions of agency/socioeconomic success, conservative–progressive beliefs, and communion, as listed in Table 1. These trait pairs were selected to capture a broad range of stereotype dimensions to distinguish groups and are supported and validated by extensive literature in social psychology (Koch et al., 2021; Abele et al., 2020).
If a group (e.g. immigrant, Asian person) has a high degree of association with a trait (e.g. religious, confident), then we consider that trait a stereotype of the group. For example, Figure 1 is the stereotype map of the group Asian people collected from our human study across the four languages that we study.
For the groups, we picked groups listed in Table 2: shared groups with shared stereotypes (groups that are present in all four countries and are expected to be targeted by similar stereotypes), shared groups with non-shared stereotypes (groups that are present in all four countries but expected to be targeted by dissimilar stereotypes), and non-shared groups (groups that exist uniquely in each country). For shared groups, we manually selected groups from the list of social groups from Cao et al. (2022) and categorized them as Shared/Shared or Shared/Non-shared. We verify the division with the human study. To collect non-shared groups, we surveyed native speakers. For each language, we asked native speakers to list social groups that they believe are unique to their culture. We then chose social groups per language based on the outcome of the majority vote.
3.2. Human stereotypes
@@
Survey Design: To collect human stereotypes, we conducted a human study on Prolific666https://www.prolific.co/ with native speakers from the United States, Russia, China, and India777Approved by our institutional IRB.. Participants were selected based on their status as either current or former residents of these countries and fluency in their respective native languages. The survey was administered in English for the U.S. and translated by native speakers into Chinese, Hindi, and Russian for the other regions. In the survey, participants first selected at least four social groups with which they were familiar. For each group, they were asked to rate associations between 16 trait pairs and the chosen social groups. We ensured a minimum of five annotations per social group in each language for both commonly shared social groups and shared groups that have unique stereotypes. For non-shared groups with unique stereotypes in a given language, we collected annotations only in that language. Each participant was compensated $2.00 for evaluating five social groups on 16 trait pairs, which took on average 10 minutes to complete. For more details about the survey design, see Appendix A.1
Annotation quality control: Annotation quality verification for subjective tasks is a challenge as there is no ground truth exists. To ensure the maximum quality of the annotators, we implemented quality testing procedures. First, we selected only participants with a higher approval rate of 90%, which we had to balance with the participants’ availability. Then we added 3 test questions to the survey to verify their engagement and consistency. For more details about questions, see Appendix A.2. A total of 286 participants completed the study, of which 151 passed the quality checks, emphasizing the importance of rigorous quality control procedures.
Participants demographics: We collected demographic data on gender, age, education, and, for non-English speakers, their consumption of American social media, with participants free to skip questions. Gender distribution was balanced across all languages (49% male, 45% female, 5% non-binary/transgender), and education levels were similar for non-English speakers (36% held bachelor’s degrees, 32% had master’s degrees, 7% held Ph. D.s). English speakers had no Ph.D. holders and a higher proportion of high school graduates (35%). Most participants were younger, with 42% aged 18-30. Russian speakers reported the highest frequency of reading American media (44%). More details on participants’ demographics are available in Appendix subsection A.3.
3.3. Model stereotypical associations
@@
To measure stereotypical group-trait associations in large language models, we adopt different approaches for different MLLMs. For mT5, we use the increased log probability score (ILPS) (Kurita et al., 2019), which computes the likelihood for the model to generate a trait given a templated sentence about a group. For example, indicates the correlation between competent and Asian people.
@@
For mBERT, we use the sensitivity test (SeT) (Cao et al., 2022), which is shown to have better alignment with human stereotypes (Cao et al., 2022). It measures how much the model weights would have to change to have the trait be the most likely generated words given a templated sentence about a group. SeT captures the model’s confidence in predicting the trait given the group.
@@
In evaluating GPT-3.5, we analyze the model in a generative context, which is its designated use case. We measure each group’s association with various traits through pairwise comparisons. We prompt the model to generate a story about the group centered around one of the traits from the trait pairs. We utilize the following prompt in English and request translations into Russian, Chinese, and Hindi from native speakers to assess the model across different languages. For each group and trait pair, we query the model ten times and compute the normalized probability score for each trait. Subsequently, we average the trait scores from all trait pair comparisons to establish the group-trait association score. To ensure equitable comparisons among positive and negative traits, we juxtapose positive traits with positive ones and negative traits with negative ones.
- system prompt::
You are a writer.
- prompt::
You are writing about {GROUP}. Before writing, think about what theme you want to pick. You can choose either ‘{TRAIT1}’ or ‘{TRAIT2}’ as your theme. You can also choose ‘neither’ if you think neither of these themes fits. Note that you can choose only one theme. Output the exact name of the theme only, without any punctuation.
When processing the outputs of GPT-3.5, we employ an exact match criterion to assign scores to traits. For traits comprising sub-tokens, we sum the log probabilities of the sub-tokens to determine the trait’s score. In instances where outputs do not precisely match the traits in the prompt, we have them manually processed by native speakers of the respective language.
Throughout this process, we consistently observe system failures or the generation of stereotypical outputs for marginalized groups. Please refer to Appendix B.1 for some examples. These occurrences can potentially cause both representational and quality of service harm to stakeholders of the model. While we do not explicitly analyze these patterns, we believe future research should investigate them thoroughly.
| Mono-BERT | EN | RU | ZH | HI |
@@
4. Stereotype Leakage and Its Effects
In this section, we present our quantitative and qualitative results of the assessment of stereotype leakage across languages in MLLMs.
Our quantitative analysis measures stereotype leakage by assessing how stereotypical associations in target language models are influenced by human stereotypes from source languages. We use mixed-effect models to quantify this leakage and identify significant cross-language influences.
Qualitatively, we examine specific stereotypical associations that leak from one language to another, exploring their nature and implications. We investigate both positive and negative stereotypes888When we say “positive stereotypes”, we mean stereotypical associations with positive traits. Note that stereotypes, regardless of how harmless they may seem, foster the essentialization of people, as well as non-polar associations, to provide a comprehensive understanding of how stereotypes are transmitted across languages.
4.1. Quantitative Results
In our analysis of mBERT, we observe significant leakages of stereotypes from Hindi to English and Chinese with coefficients of () and (), respectively. We also observe English human stereotypes manifesting in mBERT Hindi with a coefficient of (). Within the mT5 model, we find two significant stereotype leakages, both of which are leakages targeting Hindi. Russian and Chinese human stereotypes manifest in mT5 Hindi with coefficients of () and (), respectively. For GPT-3.5, we observe the most significant stereotype leakages across languages, totaling seven. We see most stereotypes leaking from English to all three other languages. The largest flows are from English to Chinese and Hindi, with coefficients of (). Meanwhile, all languages are prone to be affected by leakages from other languages, even English.